The present age of a man is four times that of his son. six years ago his age was ten times that of his son. find the present age of the man ?
Answer by TimothyLamb(4379)

m = 4s
m - 6 = 10(s - 6)
4s - 6 = 10s - 60
6s = 54
---
s = 9
m = 36
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php